OpenStudy (anonymous):

find or evaluate the integral csc2xdx

OpenStudy (anonymous):

is it -ln absv(cscx + cotx) x^2?

zepdrix (zepdrix):

\[\Large\bf\sf (\tan x)'\quad=\quad \sec^2x\]\[\Large\bf\sf (\cot x)'\quad=\quad -\csc^2 x\]

zepdrix (zepdrix):

That second derivative identity should help us out, yes? :o

OpenStudy (anonymous):

zep zep zep!

zepdrix (zepdrix):

It is batman!! :O Protect the streets good sir!

OpenStudy (anonymous):

oh but i thought the integral of csc was -ln absv(cscx + cotx)???

zepdrix (zepdrix):

Yes, when we don't have a square on it.

OpenStudy (anonymous):

there is no square its csc 2x dx

zepdrix (zepdrix):

Is your problem csc2x or csc^2x?

zepdrix (zepdrix):

oh...

OpenStudy (anonymous):

my b!

zepdrix (zepdrix):

If you need, you can do a simple u-sub to get things to work out for you. \[\Large\bf\sf u=2x, \qquad\qquad \frac{1}{2}du=dx\] \[\Large\bf\sf \int\limits \csc 2x\;dx\quad=\quad \frac{1}{2}\int \csc u\; du \]Now you can apply that rule directly to the csc u

OpenStudy (anonymous):

oh so just 1/2 csc (2x)(2)

zepdrix (zepdrix):

No you didn't integrate yet :o Integrate the csc u

zepdrix (zepdrix):

confused?

zepdrix (zepdrix):

You just posted the integral of csc a minute ago .. :P

OpenStudy (anonymous):

OH OOPS! right so take the integral and since its a negative ln can i just move that in front of the 1/2

OpenStudy (anonymous):

i.e. -1/2 ln absv (cscx+cotx)

zepdrix (zepdrix):

\[\Large\bf\sf =\quad -\frac{1}{2}\ln\left|\csc u+\cot u\right|+c\]Mmmm yah looks good so far :)

zepdrix (zepdrix):

We're integrating in u, yes? So our stuff inside the trig functions should be u's.

OpenStudy (anonymous):

True! and then you can substitute the u( 2x) back in?

zepdrix (zepdrix):

ya there we go!

OpenStudy (anonymous):

and i forget the c is substituted with du so instead of plus c it would be plus 2?

zepdrix (zepdrix):

c is substituted with du...? :o

OpenStudy (anonymous):

or do you just get rid of the c altogether???

OpenStudy (anonymous):

constant?

OpenStudy (anonymous):

indefinite integral, so you add +C if that's what you're asking o_O

zepdrix (zepdrix):

c stays. We made a substitution, Integrated, (an arbitrary constant of integration shows up) undo our substitution, yah the c has nothing to do with changing from x to u. It's still arbitrary regardless of our variable we're integrating in :o

OpenStudy (anonymous):

ok ok! and does the problem have to be simplified now or is it done???

OpenStudy (anonymous):

bc it looks rlly messy...

OpenStudy (anonymous):

Leave it

zepdrix (zepdrix):

Yah it sure looks messy :) But it can't be written much simpler than that. That's just how some of these work out.

OpenStudy (anonymous):

^^^^

OpenStudy (anonymous):

LOVELY! THANKS AGAIN:-)